A bank makes a 30 year Fully Amortizing FRM for $1,000,000 at an annual interest rate of 4.13% compounded monthly, with monthly payments. What is the market value of this loan after 7 years of payments if the annual interest rate for this loan is 10% compounded monthly?"
Answer #1

Loan monthly payment=1000000*(4.13%/12)/(1-1/(1+4.13%/12)^(12*30))=4849.402724

Market value of loan=4849.402724/(4.13%/12)*((1+4.13%/12)^(12*(30-7))-1)
=2228007.681